OpenStudy (anonymous):

please help Simplify square root of 2 multiplied by the cube root of 2

OpenStudy (solomonzelman):

\[\sqrt{2} \times \sqrt[3]{2}\]

OpenStudy (solomonzelman):

this is also same as saying \[\sqrt[2]{2} \times \sqrt[3]{2}\]

OpenStudy (solomonzelman):

the formula for this is as follows.\[\sqrt[c]{a^b}=a^{b/c}\]

OpenStudy (solomonzelman):

\[2^{1/2} \times 2^{1/3}\]

OpenStudy (solomonzelman):

\[2^{1/2+1/3}\]

OpenStudy (solomonzelman):

\[2^{5/6}\]

OpenStudy (solomonzelman):

\[\sqrt[6]{2^5}\]

OpenStudy (anonymous):

alright thanks

OpenStudy (solomonzelman):

Anytime!